A rectangle has an area of 144cm^2 and a perimeter of 68cm. What are the lengths of the sides?
Answer by stanbon(75887)

A rectangle has an area of 144cm^2 and a perimeter of 68cm. What are the lengths of the sides?
------
Equations:
LW = 144
2(L+W) = 68
-----
LW = 144
L + W = 34
-----
L = 34-W
-------
Substitute and solve for "W":
(34-W)W = 144
-w^2 + 34w - 144 = 0
w^2 - 34w + 144 = 0
---
w = [34 +- sqrt(34^2 -4*144)]/2
----
w = [34 +- sqrt(580)]/2
Positive solution:
w = (17 + sqrt(145))
W is appoximately 29.04 cm
----
Since L + W = 34, L = 34-29.04 = 4.96 cm
======================================================
Cheers,
stan H.